OpenStudy (anonymous):

Use completing the square to describe the graph of the following function, support your answer graphically. f(x)=-2x^2+4x+5

OpenStudy (gorv):

f=-2(x^2+2x+5/2) =-2(x^2+2x+1+3/2) =-2(x^2+2x+1)-2*3/2 =-2(x+1)^2-3

OpenStudy (anonymous):

@greenlegodude57 can you help? im not sure what that guy did =(

OpenStudy (anonymous):

I'm not sure about this one, sorry. @gorv Could you explain to her?

OpenStudy (anonymous):

ok thanks @greenlegodude57

OpenStudy (gorv):

we need to make perfect square @jdorta1

OpenStudy (gorv):

r u there ??

OpenStudy (anonymous):

im horrible at this stuff, so if you can explain how to do it =(

OpenStudy (anonymous):

yes

OpenStudy (gorv):

x^2 is there so make a a perfect square using this

OpenStudy (gorv):

x^2+2*a*x+a^2=(x+a)^2

OpenStudy (gorv):

take -2 c0mmon

OpenStudy (anonymous):

I'm bad at the "perfect square" thing too.

OpenStudy (anonymous):

ok

OpenStudy (radar):

The problem wants you to find the zeroes by "completing the square" f(x)= -2x^2 + 4x +5 = 0 First step get the coefficient of the "x^2" term to equal a 1. To do this we divide the equation by -2 (both sides) x^2 - 2x -5/2 = 0 (0 divided by -2 is still 0) Next step as 5/2 to both sides in order to get the constant on the right hand side. x^2 -2x = 0 +5/2 Now divide the coefficient of the x term by 2 (taking half of it), then square it and add it to both sides of the equation x^2 -2x + 1 = 5/2 + 1. Now on the left side you have the perfect square, express it so. (x - 1)^2 = 7/2 (adding the right hand side. Do you follow the steps so far? Because the finale is near.

OpenStudy (radar):

The final step is to take the square root of both sides and solve for x.

OpenStudy (radar):

\[\sqrt{(x - 1)2}=\sqrt{7/2}\]\[x - 1 =\sqrt{7/2}\]\[x = 1\pm \sqrt{7/2}\]The radical can be further simplified. Do you know how to "rationalize" the denominator?

OpenStudy (radar):